# Cold day-side winds shape large leading streams in evaporating exoplanet atmospheres

### v1

@@ 3.1 Formation of a predominantly leading stream
In addition to investigating the orbital mid-plane, we also examined the vertical structure of the planetary wind. Fig. 4 presents the helium equivalent width light curve along with the surface density of the planetary wind throughout the planetary orbit of the fiducial model S2. We show one light curve resulting from a density scaling factor of , and also include results from a higher scaling factor of . While the latter overestimates the in-transit absorption, it more accurately matches the tail absorption. A possible explanation for why our models cannot fully capture the density gradient between the tails and the planet is that they do not account for thermodynamics, cooling or heating rates.
@@
Looking at the vertical structure, the outflow from the day side of the planet converges at the phase , covering a smaller part of the stellar disk and causing a dip in the light curve. As the phase progresses earlier, the outflow disperses into irregular patches and covers a larger part of the stellar disk, which explains the local maximum in the light curve at . The density distribution is highly variable due to instabilities, which is reflected in both the observed light curve and our model. In our model, the orbit-to-orbit variability is about 20 m in equivalent width, while the observed spread can reach up to m (e.g., at for HAT-P-67 b).
@@ 3.2 Characteristics of the helium line profile caused by the stream morphology
In Fig. 5, we present the spectra of the fiducial model S2 in sequential order, progressing from pre-transit (blue) to mid-transit (gray) and post-transit (red). The line profiles show pronounced irregularities, deviating significantly from a thermally broadened Gaussian profile. The mid-transit line profile reveals three distinct features, each marked with corresponding numbers. Generally, in addition to the He i 1083 nm feature being a triplet, we would expect a kinematically broadened line showing three peaks, each corresponding to distinct velocity regimes of the planetary wind. The first, blueshifted peak is formed by gas launched from the night side that moves toward the observer. The third peak is redshifted, originating from the day-side wind moving away from the observer. Additionally, there is a line component with little to no shift, which forms in a high-density region close to the planet’s surface, resulting in the second peak.
@@
This is what we observe in Fig. 5; however, the formation of the first peak is actually more complex. It is not just a part of the helium main component (the unresolved lines of and ) that is blueshifted as is forms in the gas launched from the night side; the blue component () of the third peak also contributes to the first peak. As the helium line forming in a high-density region is optically thick, the blue component is significantly enhanced. Consequently, the first peak is a combination of the blueshifted main component and the redshifted blue component from the third peak.
@@
The orbit-to-orbit variability, as indicated by the standard deviation calculated from the average of five snapshots, is low, while the three aforementioned features persist. It is also notable that the third peak of the mid-transit spectrum shows the highest absorption. This is due to the day night side anisotropy, with higher gas density on the day side of the planet. If the features two and three are observationally not resolved, the overall line may appear redshifted, while the exact position and strength of the three peaks depend on the density and velocity distribution of the specific planetary wind. Notably, the shape of the highly resolved He i 1083 nm line observed in HAT-P-32 b by Czesla et al. (2022) (see their Fig. 18) shows similar irregularities, indicating a kinematically broadened line. While instrumental uncertainties cannot be entirely ruled out, the blue component of the triplet appears to be double-peaked, and on the longer wavelength side of the main component, two smaller peaks are visible. The small ratio between the blue and main components further suggests that the line is optically thick.
@@
For our fiducial model S2, the absorption after the transit (red spectra) is only significant up to the phase . At higher orbital phases, the absorption diminishes, aligning with the sharp edge in the light curve (see Fig. 3). In the snapshot of simulation S2 (Fig. 2, bottom right) a trailing tail is visible, however, the gas density probed at higher phase angles is not high enough to contribute to the line formation. As illustrated in Fig. 3, pre-transit absorption extends up to in orbital phase. In contrast to the in-transit spectra, during the pre-transit phase (), both the line profile and equivalent width display high variability. This indicates that the leading arm is exposed to instabilities from the interaction with the stellar wind. The snapshot in Fig. 2 indeed shows Kelvin-Helmholtz and Rayleigh-Taylor instabilities at phases .
3.3 Wind-wind interactions influence on tails
@@
Furthermore, we found that the velocity shift of the metastable helium line provides valuable insights into constraining the strength of the stellar wind. The left panel in Fig. 6 illustrates the velocity shift of the helium line across different stellar wind scenarios throughout the orbital phase. Given the highly irregular line profiles, we opted to calculate the weighted mean of the spectra to determine the velocity shift.
In the case of a weak stellar wind (model S2, blue), the velocity shifts during mid-transit follow the orbital movement of the planet. During ingress, the helium line is redshifted relative to the planetary movement, while during egress it is blueshifted. These shifts suggest that, during ingress, the probed gas in the leading stream is accelerated away from the planet towards the star (towards the point). Conversely, during egress, the gas in the trailing stream is accelerated away from the planet towards the observer (towards the point).
After passing through the Lagrangian points, the gas forms tidal streams that orbit the host star in a Keplerian manner, with low radial velocities compared to the planet’s velocity. The absorption signatures of these streams therefore remain close to the stellar rest frame. In theory, the gas in the leading stream moves with a higher angular velocity as it orbits the star at a shorter distance. Similarly, the gas in the trailing stream moves slightly slower, orbiting at a greater distance.
With increasing strength of the stellar wind, both pre- and post-transit spectra show progressively greater blueshifts. In contrast to S2, the velocity shifts in the transit of the stronger stellar wind models S3 and S4 no longer follow the planetary motion. As the metastable helium densities around the planet are so high in these scenarios, the line forms predominantly in the turbulent region further away from the planet. This gas is pushed towards the observer by the stellar wind, similar to the tails, resulting in a blueshift of the spectral line.
@@
This analysis emphasizes the significance of two key measurements to constrain the planetary outflow with our simulations. Both the equivalent width and the velocity shift over orbital phase are important. The light curve reveals the extent of the outflow, while the velocity shift provides insights into its kinematics. Importantly, the velocity shift is intricately linked, as demonstrated, to the strength of the stellar wind. However, the light curve serves to rule out a stellar wind scenario that is too strong, thereby establishing it as the more robust analysis tool.
@@ 4 Discussion
In comparison to our previous work Nail et al. (2024), which explored the effects of anisotropic outflows in a hot wind scenario, this study focuses on the behavior of a cold planetary wind. In Nail et al. (2024), a significantly enhanced day side outflow resulted in a more extended cavity around the planet before the gas was shocked and pushed behind the planet by the stellar wind. Additionally, a correlation was observed between the blueshift of the metastable helium line and the degree of anisotropy. In contrast, for the cold planetary wind studied here, the spectral lines are broadened kinematically. During transit, this broadening is primarily driven by the two streams diverging in opposite directions, whereas outside of transit, the line shift is mainly influenced by the strength of the stellar wind.
The question arises as to why the planetary winds of HAT-P-67 b and HAT-P-32 b should be cold and not hot. A contributing factor for a cooler outflow is that the effective potential for overcoming the Roche lobe is relatively low, since the planets are close to their host stars while having a low gravitational potential. The gas can escape more easily from the Roche lobe and is not further heated due to the expansion cooling.
The second question that arises is why the outflow is asymmetric, with a more pronounced leading stream. While we remain agnostic about the wind-driving mechanisms in this study, it becomes clear that a day-side dominated outflow can reproduce an asymmetric light curve. This suggests that stellar radiation indeed plays a role in driving the escape. Further supporting this theory is the observation that the outflow of HAT-P-67 b has apparently a greater degree of asymmetry than HAT-P-32 b. HAT-P-32 is a main sequence star, while HAT-P-67 has recently undergone reinflation and is either at the terminal stage of the main sequence or has evolved into a sub-giant (Gully-Santiago et al., 2023). Consequently, HAT-P-67 b currently receives approximately twice the incident flux compared to what it would have received at the Zero Age Main Sequence (Zhou et al., 2017). This enhanced irradiation may enable a stronger atmospheric escape from the planet’s day side compared to HAT-P-32 b.
@@ 5 Conclusion
We have demonstrated through our models that the characteristics of a leading outflow stream as observed through He i 1083 nm observations in HAT-P-32 b and HAT-P-67 b (see Fig. 1) are best explained by a high planetary mass-loss rate of in the order of to g s-1 ( yr-1) and a relatively cool outflow temperature of approximately K, or just a few times the equilibrium temperature, primarily originating from the heated day side of the planet.
Our models suggest that to create a stream with the majority of absorption occurring pre-transit, a relatively cool outflow is essential (see Fig. 2). A day-side dominated outflow is favored over an isotropic outflow, as it shows greater asymmetry in the light curve (see Fig. 3).In contrast, the hot wind scenarios fail to replicate the extensive light curve of excess He i 1083 nm absorption across orbital phase.
The He i 1083 nm line of the cold wind scenario shows a large absorption depth and an irregular line profile. Our analysis reveals that the line is not broadened thermally, but instead kinematically. The irregular shape arises from distinct velocity distributions within the outflow (Fig. 5). By measuring both the equivalent width and velocity shift over orbital phase, it is possible to constrain both the outflow and its shaping by the stellar wind (Fig. 3 and Fig. 6).
@@ Acknowledgements.
- Allart et al. (2018) Allart, R., Bourrier, V., Lovis, C., et al. 2018, Science, 362, 1384, _eprint: https://www.science.org/doi/pdf/10.1126/science.aat5879
- Bello-Arufe et al. (2023) Bello-Arufe, A., Knutson, H. A., Mendonça, J. M., et al. 2023, The Astronomical Journal, 166, 69
@@
- Carolan et al. (2021) Carolan, S., Vidotto, A. A., Hazra, G., D’Angelo, C. V., & Kubyshkina, D. 2021, Monthly Notices of the Royal Astronomical Society, 508, 6001, arXiv: 2110.05200
- Carroll-Nellenback et al. (2017) Carroll-Nellenback, J., Frank, A., Liu, B., et al. 2017, Monthly Notices of the Royal Astronomical Society, 466, 2458, arXiv: 1604.08532
- Czesla et al. (2024) Czesla, S., Lampón, M., Cont, D., et al. 2024, Astronomy & Astrophysics, 683, A67
- Czesla et al. (2022) Czesla, S., Lampón, M., Sanz-Forcada, J., et al. 2022, Astronomy & Astrophysics, 657, A6
@@
- Fossati et al. (2010) Fossati, L., Bagnulo, S., Elmasli, A., et al. 2010, The Astrophysical Journal, 720, 872
- Guilluy et al. (2020) Guilluy, G., Andretta, V., Borsa, F., et al. 2020, Astronomy & Astrophysics, 639, A49
- Gully-Santiago et al. (2023) Gully-Santiago, M., Morley, C. V., Luna, J., et al. 2023, A Large and Variable Leading Tail of Helium in a Hot Saturn Undergoing Runaway Inflation, arXiv:2307.08959 [astro-ph]
- Hartman et al. (2011) Hartman, J. D., Bakos, G. A., Torres, G., et al. 2011, The Astrophysical Journal, 742, 59
@@
- Haswell et al. (2012) Haswell, C. A., Fossati, L., Ayres, T., et al. 2012, The Astrophysical Journal, 760, 79
- Jensen et al. (2018) Jensen, A. G., Cauley, P. W., Redfield, S., Cochran, W. D., & Endl, M. 2018, The Astronomical Journal, 156, 154
- Kreidberg & Oklopčić (2018) Kreidberg, L. & Oklopčić, A. 2018, Research Notes of the AAS, 2, 44
- Lampón et al. (2023) Lampón, M., López-Puertas, M., Sanz-Forcada, J., et al. 2023, Astronomy & Astrophysics, 673, A140
- Llama et al. (2011) Llama, J., Wood, K., Jardine, M., et al. 2011, Monthly Notices of the Royal Astronomical Society: Letters, 416, L41
- MacLeod & Oklopčić (2022) MacLeod, M. & Oklopčić, A. 2022, The Astrophysical Journal, 926, 226
- McCann et al. (2019) McCann, J., Murray-Clay, R. A., Kratter, K., & Krumholz, M. R. 2019, The Astrophysical Journal, 873, 89
@@
- Nichols et al. (2015) Nichols, J. D., Wynn, G. A., Goad, M., et al. 2015, The Astrophysical Journal, 803, 9
@@
- Oklopčić & Hirata (2018) Oklopčić, A. & Hirata, C. M. 2018, The Astrophysical Journal, 855, L11, arXiv: 1711.05269
- Owen (2019) Owen, J. E. 2019, Annual Review of Earth and Planetary Sciences, 47, 67, arXiv: 1807.07609
- Salz et al. (2018) Salz, M., Czesla, S., Schneider, P. C., et al. 2018, Astronomy & Astrophysics, 620, A97
- Sicilia et al. (2024) Sicilia, D., Scandariato, G., Guilluy, G., et al. 2024 [arXiv:2404.03317]
- Spake et al. (2021) Spake, J. J., Oklopčić, A., & Hillenbrand, L. A. 2021, The Astronomical Journal, 162, 284
@@
- Stone et al. (2020) Stone, J. M., Tomida, K., White, C. J., & Felker, K. G. 2020, The Astrophysical Journal Supplement Series, 249, 4
@@
- Vidotto & Cleary (2020) Vidotto, A. A. & Cleary, A. 2020, Monthly Notices of the Royal Astronomical Society, 494, 2417
- Wang et al. (2019) Wang, Y.-H., Wang, S., Hinse, T. C., et al. 2019, The Astronomical Journal, 157, 82
- Yan et al. (2024) Yan, D., Guo, J., Seon, K.-i., et al. 2024, Astronomy & Astrophysics, 686, A208
- Zhang et al. (2023) Zhang, M., Knutson, H. A., Dai, F., et al. 2023, Detection of Atmospheric Escape from Four Young Mini Neptunes, arXiv:2207.13099 [astro-ph]
- Zhou et al. (2017) Zhou, G., Bakos, G. A., Hartman, J. D., et al. 2017, The Astronomical Journal, 153, 211
Appendix A Additional material

### v2

@@ 3.1 Formation of a predominantly leading stream
In addition to investigating the orbital midplane, we also examined the vertical structure of the planetary wind. Figure 4 presents the helium EW light curve along with the surface density of the planetary wind throughout the planetary orbit of the fiducial model S2. We show one light curve resulting from a density scaling factor of and also include results from a higher scaling factor of . While the latter overestimates the in-transit absorption, it more accurately matches the tail absorption. A possible explanation for why our models cannot fully capture the density gradient between the tails and the planet is that they do not account for thermodynamics, cooling or heating rates.
@@
Looking at the vertical structure, the outflow from the dayside of the planet converges at the phase due to tidal compression, covering a smaller part of the stellar disk and causing a dip in the light curve. As the phase progresses earlier, the outflow disperses into irregular patches and covers a larger part of the stellar disk, which explains the local maximum in the light curve at . The density distribution is highly variable due to instabilities, which is reflected in both the observed light curve and our model. In our model, the orbit-to-orbit variability is about 20 m in EW, while the observed spread can reach up to m (e.g., at for HAT-P-67 b).
@@ 3.2 Characteristics of the helium line profile caused by the stream morphology
In Fig. 5 we present the spectra of the fiducial model S2 in sequential order, progressing from pre-transit (blue) to mid-transit (gray) and post-transit (red). The line profiles show pronounced irregularities, deviating significantly from a thermally broadened Gaussian profile. The mid-transit line profile reveals three distinct features, each marked with corresponding numbers. Based on the velocity distribution observed in the snapshots of the stream morphology (see Fig. 8), we expect the He i 1083 nm feature to show kinematic broadening, with three distinct peaks. These peaks would correspond to different velocity regimes within the planetary wind, in addition to the He i 1083 nm feature being a triplet. The first, blueshifted peak is formed by gas launched from the nightside that moves toward the observer. The third peak is redshifted, originating from the dayside wind moving away from the observer. Additionally, there is a line component with little to no shift, which forms in a high-density region close to the planet’s surface, resulting in the second peak.
@@
This is what we observe in Fig. 5; however, the formation of the first peak is actually more complex. It is not just a part of the helium main component (the unresolved lines of and ) that is blueshifted as is forms in the gas launched from the nightside; the blue component () of the third peak also contributes to the first peak. As the helium line forming in a high-density region is optically thick, the blue component is significantly enhanced. Consequently, the first peak is a combination of the blueshifted main component and the redshifted blue component from the third peak. In Appendix B we analyze the line-forming regions, focusing exclusively on the singlet line to isolate the contributions of the triplet components and the three kinematic peaks.
@@
The orbit-to-orbit variability, as indicated by the standard deviation calculated from the average of five snapshots, is low, while the three aforementioned features persist. It is also notable that the third peak of the mid-transit spectrum shows the highest absorption. This is due to the day nightside anisotropy, with higher gas density on the dayside of the planet. If features 2 and 3 are not observationally resolved, the overall line may appear redshifted, while the exact position and strength of the three peaks depend on the density and velocity distribution of the specific planetary wind. Notably, the shape of the highly resolved He i 1083 nm line observed in HAT-P-32 b by Czesla et al. (2022, see their Fig. 18) shows similar irregularities, indicating a kinematically broadened line. While instrumental uncertainties cannot be entirely ruled out, the blue component of the triplet appears to be double-peaked, and on the longer wavelength side of the main component, two smaller peaks are visible. The small ratio between the blue and main components further suggests that the line is optically thick.
@@
For our fiducial model S2, the absorption after the transit (red spectra) is only significant up to the phase . At higher orbital phases, the absorption diminishes, aligning with the sharp edge in the light curve (see Fig. 3). In the snapshot of simulation S2 (Fig. 2, bottom right) a trailing tail is visible; however, the gas density probed at higher phase angles is not high enough to contribute to the line formation. As illustrated in Fig. 3, pre-transit absorption extends up to in orbital phase. In contrast to the in-transit spectra, during the pre-transit phase (), both the line profile and EW display high variability. This indicates that the leading arm is exposed to instabilities from the interaction with the stellar wind. The snapshot in Fig. 2 indeed shows Kelvin-Helmholtz and Rayleigh-Taylor instabilities at phases .
Furthermore, we explored the line profiles for model S1, the isotropic stream scenario. In this case, the three peaks during transit are also visible, with the key difference being that the post-transit spectra show higher absorption depths. However, the overall shapes of the lines remain unchanged. In the isotropic case (S1), the overall spectrum appears less redshifted compared to the anisotropic case (S2). This is because the cells contributing to line formation are more symmetrically distributed around the planet in model S1. In contrast, in the anisotropic case (S2), the cells are predominantly located in the leading stream, where higher densities occur due to the day-night anisotropy (see Fig. 8).
3.3 Wind–wind interactions’ influence on tails
@@
Furthermore, we find that the velocity shift of the metastable helium line provides valuable insights into constraining the strength of the stellar wind. The left panel in Fig. 6 illustrates the velocity shift of the helium line across different stellar wind scenarios throughout the orbital phase. Given the highly irregular line profiles, we opted to calculate the weighted mean of the spectra to determine the velocity shift. Each wavelength is assigned a weight based on its corresponding absorption depth. Using this weighting, a weighted average wavelength is calculated, representing the central position of the helium spectrum. This wavelength, representing the central position of the helium spectrum, is then converted into a velocity shift relative to the centroid of the helium triplet’s main component.
In the case of a weak stellar wind (model S2, blue), the velocity shifts during mid-transit follow the orbital movement of the planet. During ingress, the helium line is redshifted relative to the planetary movement, while during egress it is blueshifted. These shifts suggest that, during ingress, the probed gas in the leading stream is accelerated away from the planet toward the star (toward the point). Conversely, during egress, the gas in the trailing stream is accelerated away from the planet toward the observer (toward the point).
After passing through the Lagrangian points, the gas forms tidal streams that orbit the host star in a Keplerian manner, with low radial velocities compared to the planet’s velocity. The absorption signatures of these streams therefore remain close to the stellar rest frame. In theory, the gas in the leading stream moves with a higher angular velocity as it orbits the star at a shorter distance. Similarly, the gas in the trailing stream moves slightly more slowly, orbiting at a greater distance.
With increasing strength of the stellar wind, both pre- and post-transit spectra show progressively greater blueshifts. In contrast to S2, the velocity shifts in the transit of the stronger stellar wind models S3 and S4 no longer follow the planetary motion. Due to the high ram pressure of the stellar wind, the planetary outflow is strongly confined. As a result, the metastable helium densities around the planet become so high that the line primarily forms in the turbulent region farther from the planet, closer to the star (see the optical depth contours in Fig. 6, bottom right). This gas is pushed toward the observer by the stellar wind, similar to the tails, resulting in a blueshift of the spectral line.
@@
This analysis emphasizes the significance of two key measurements to constrain the planetary outflow with our simulations. Both the EW and the velocity shift over orbital phase are important. The light curve reveals the extent of the outflow, while the velocity shift provides insights into its kinematics. Importantly, the velocity shift is intricately linked, as demonstrated, to the strength of the stellar wind. However, the light curve serves to rule out a stellar wind scenario that is too strong, thereby establishing it as the more robust analysis tool.
@@ 4 Discussion
In comparison to our previous work (Nail et al., 2024), which explored the effects of anisotropic outflows in a hot wind scenario, this study focuses on the behavior of a cold planetary wind. In Nail et al. (2024), a significantly enhanced dayside outflow resulted in a more extended cavity around the planet before the gas was shocked and pushed behind the planet by the stellar wind. Additionally, a correlation was observed between the blueshift of the metastable helium line and the degree of anisotropy. In contrast, for the cold planetary wind studied here, the spectral lines are broadened kinematically. During transit, this broadening is primarily driven by the two streams diverging in opposite directions, whereas outside of transit, the line shift is mainly influenced by the strength of the stellar wind.
The question arises as to why the planetary winds of HAT-P-67 b and HAT-P-32 b should be cold, with temperatures of only a few thousand kelvins, rather than hot, reaching around 10000 K A contributing factor for a cooler outflow is that the effective potential for overcoming the Roche lobe is relatively low, since the planets are close to their host stars while having a low gravitational potential. The gas can escape more easily from the Roche lobe and is not further heated due to the expansion cooling.
The second question that arises is why the outflow is asymmetric, with a more pronounced leading stream. While we remain agnostic about the wind-driving mechanisms in this study, it becomes clear that a dayside-dominated outflow can reproduce an asymmetric light curve. This suggests that stellar radiation indeed plays a role in driving the escape. Further supporting this theory is the observation that the outflow of HAT-P-67 b has apparently a greater degree of asymmetry than HAT-P-32 b. HAT-P-32 is a main sequence star, while HAT-P-67 has recently undergone reinflation and is either at the terminal stage of the main sequence or has evolved into a subgiant (Gully-Santiago et al., 2024). Consequently, HAT-P-67 b currently receives approximately twice the incident flux compared to what it would have received at the zero age main sequence (Zhou et al., 2017). This enhanced irradiation may enable a stronger atmospheric escape from the planet’s dayside compared to HAT-P-32 b.
@@ 5 Conclusion
We have demonstrated through our models that the characteristics of a leading outflow stream as observed through He i 1083 nm observations in HAT-P-32 b and HAT-P-67 b (see Fig. 1) are best explained by a high planetary mass-loss rate on the order of to g s-1 ( yr-1) and a relatively cool outflow temperature of approximately K, which is just a few times the equilibrium temperature, primarily originating from the heated dayside of the planet.
Our models suggest that to create a stream with the majority of absorption occurring pre-transit, a relatively cool outflow is essential (see Fig. 2). A dayside-dominated outflow is favored over an isotropic outflow, as it shows greater asymmetry in the light curve (see Fig. 3). In contrast, the hot wind scenarios fail to replicate the extensive light curve of excess He i 1083 nm absorption across the orbital phase.
The He i 1083 nm line of the cold wind scenario shows a large absorption depth and an irregular line profile. Our analysis reveals that the line is not broadened thermally, but instead kinematically. The irregular shape arises from distinct velocity distributions within the outflow (Fig. 5). By measuring both the EW and velocity shift over the orbital phase, it is possible to constrain both the outflow and its shaping by the stellar wind (Figs. 3 and 6).
Data availability
We have uploaded the updated radiative transfer code together with simulation material on Zenodo (10.5281/zenodo.13988501).
@@ Acknowledgements.
- Allart et al. (2018) Allart, R., Bourrier, V., Lovis, C., et al. 2018, Science, 362, 1384
- Bello-Arufe et al. (2023) Bello-Arufe, A., Knutson, H. A., Mendonça, J. M., et al. 2023, ApJ, 166, 69
@@
- Carolan et al. (2021) Carolan, S., Vidotto, A. A., Hazra, G., D’Angelo, C. V., & Kubyshkina, D. 2021, MNRAS, 508, 6001
- Carroll-Nellenback et al. (2017) Carroll-Nellenback, J., Frank, A., Liu, B., et al. 2017, MNRAS, 466, 2458
- Czesla et al. (2024) Czesla, S., Lampón, M., Cont, D., et al. 2024, A&A, 683, A67
- Czesla et al. (2022) Czesla, S., Lampón, M., Sanz-Forcada, J., et al. 2022, A&A, 657, A6
@@
- Fossati et al. (2010) Fossati, L., Bagnulo, S., Elmasli, A., et al. 2010, ApJ, 720, 872
- Guilluy et al. (2020) Guilluy, G., Andretta, V., Borsa, F., et al. 2020, A&A, 639, A49
- Gully-Santiago et al. (2024) Gully-Santiago, M., Morley, C. V., Luna, J., et al. 2024, ApJ, 167, 142
- Hartman et al. (2011) Hartman, J. D., Bakos, G. A., Torres, G., et al. 2011, ApJ, 742, 59
@@
- Haswell et al. (2012) Haswell, C. A., Fossati, L., Ayres, T., et al. 2012, ApJ, 760, 79
- Jensen et al. (2018) Jensen, A. G., Cauley, P. W., Redfield, S., Cochran, W. D., & Endl, M. 2018, ApJ, 156, 154
- Kreidberg & Oklopčić (2018) Kreidberg, L. & Oklopčić, A. 2018, RNAAS, 2, 44
- Lampón et al. (2023) Lampón, M., López-Puertas, M., Sanz-Forcada, J., et al. 2023, A&A, 673, A140
- Llama et al. (2011) Llama, J., Wood, K., Jardine, M., et al. 2011, MNRAS Letters, 416, L41
- MacLeod & Oklopčić (2022) MacLeod, M. & Oklopčić, A. 2022, ApJ, 926, 226
- McCann et al. (2019) McCann, J., Murray-Clay, R. A., Kratter, K., & Krumholz, M. R. 2019, ApJ, 873, 89
@@
- Nichols et al. (2015) Nichols, J. D., Wynn, G. A., Goad, M., et al. 2015, ApJ, 803, 9
@@
- Oklopčić & Hirata (2018) Oklopčić, A. & Hirata, C. M. 2018, ApJ, 855, L11
- Owen (2019) Owen, J. E. 2019, Annu. Rev. Earth Planet. Sci., 47, 67
- Salz et al. (2018) Salz, M., Czesla, S., Schneider, P. C., et al. 2018, A&A, 620, A97
- Sicilia et al. (2024) Sicilia, D., Scandariato, G., Guilluy, G., et al. 2024, A&A, 687, A143
- Spake et al. (2021) Spake, J. J., Oklopčić, A., & Hillenbrand, L. A. 2021, ApJ, 162, 284
@@
- Stone et al. (2020) Stone, J. M., Tomida, K., White, C. J., & Felker, K. G. 2020, ApJS, 249, 4
@@
- Vidotto & Cleary (2020) Vidotto, A. A. & Cleary, A. 2020, MNRAS, 494, 2417
- Wang et al. (2019) Wang, Y.-H., Wang, S., Hinse, T. C., et al. 2019, ApJ, 157, 82
- Yan et al. (2024) Yan, D., Guo, J., Seon, K.-i., et al. 2024, A&A, 686, A208
- Zhang et al. (2023) Zhang, Z., Morley, C. V., Gully-Santiago, M., et al. 2023, Science Advances, 9, eadf8736
- Zhou et al. (2017) Zhou, G., Bakos, G. A., Hartman, J. D., et al. 2017, ApJ, 153, 211
Appendix A Simulation movie